# Photon transport and blockade based on non-Markovian interactions between a microring resonator and waveguide

### v1

@@ V Conclusion
In summary, we have delved into the photon transport and blockade, particularly focusing on the non-Markovian interactions between a microring resonator and a waveguide. By employing the two-point coupling method, we have uncovered how a time delay can be induced for the propagating photon within the waveguide, subsequently impacting both single-photon and two-photon transports. As a result, by tuning the waveguide length between two coupled points, the two input photons cannot be ideally transmitted, leading to the manifestation of photon blockade. Furthermore, we have also explored the generation of photon states within the resonator by applying a classical drive upon the resonator. Through this approach, we have discerned that the occurrence of photon blockade, as well as the modulation of photon number states, can be achieved by tuning the delayed phase in the non-Markovian system. Our research explores the intricate dynamics of photon transport and blockade in the framework of non-Markovian interactions between a microring resonator and a waveguide. This exploration presents intriguing possibilities for further exploration and application in the realm of quantum optics, quantum networking, and room-temperature quantum information processing.
@@ Appendix A Derivation on non-Markovian single-photon transport
Then we have the following Fourier transforms for $\mathcal{A}(t)$ and $\dot{\mathcal{A}}(t)$ as
@@ Appendix B Derivation on non-Markovian two-photon transport
In combination with Eq. (14a) and the Fourier inversion of Eqs. (22,23)
@@
Due to the mathematical structure of the two-photon Hilbert space [32], the scattering matrix for two-photon transport can be derive as Eq. (27) in the main text.

### v2

@@ V Conclusion
In summary, we have delved into the photon transport and blockade, particularly focusing on the non-Markovian interactions between a microring resonator and a waveguide. By employing the two-point coupling approach from the perspective of scattering matrices, we have uncovered how a time delay can be induced by the propagating photons in the waveguide, subsequently impacting both single-photon and two-photon transports. As a result, by tuning the waveguide length between two coupled points, the two input photons cannot be ideally transmitted, leading to the manifestation of photon blockade. Furthermore, we have also explored the generation of photon states within the resonator by applying a classical drive upon the resonator. Through this approach, we have discerned that the occurrence of photon blockade, as well as the modulation of photon number states, can be achieved by tuning the delayed phase in the non-Markovian system. Our research explores the intricate dynamics of photon transport and blockade in the framework of non-Markovian interactions between a microring resonator and a waveguide. This exploration presents intriguing possibilities for further exploration and application in the realm of quantum optics, quantum networking, and room-temperature quantum information processing.
@@ Appendix A Derivation on non-Markovian single-photon transport
Based on the following Fourier transforms for $\mathcal{A}(t)$ and $\dot{\mathcal{A}}(t)$,
@@ Appendix B Derivation on non-Markovian two-photon transport
In combination with Eq. (14a) and the Fourier inversions of Eqs. (22,23)
@@
Due to the mathematical structure of the two-photon Hilbert space [59], the scattering matrix for two-photon transport can be derived as Eq. (27) in the main text.